OpenStudy (anonymous):

(x)^2+(1/2x+11)^2=(2x+1)^2 PLEASE HELP

OpenStudy (anonymous):

\[(x)^2+(\frac{ 1 }{ 2 }x+11)^2=(2x+1)^2\]

OpenStudy (mathmate):

Expand the expressions to get \( -\frac{11}{4}x^2+7x+120=0 \) Solve for x.

OpenStudy (anonymous):

x^2= -480/11

OpenStudy (mathmate):

I have a negative fraction and a positive integer as the two possible answers.

OpenStudy (anonymous):

how did you get that @mathmate

OpenStudy (anonymous):

\[x ^{2}+\frac{ 1 }{ 4}x ^{2}+2*\frac{ 1 }{ 2}x*11+121=4x ^{2}+2*2x*1+1^{2}\] \[x ^{2}+\frac{ 1 }{ 4 }x ^{2}+11x+121-4x ^{2}-4x-1=0\] \[-3x ^{2}+\frac{ 1 }{ 4 }x ^{2}+7x+120=0\] \[-12x ^{2}+x ^{2}+28x+120=0\] \[11x ^{2}-28x-120=0\] \[x=\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\] a=11, b=-28,c=-120 calculate for x

OpenStudy (mathmate):

@surjithayer I believe c should be -480. You probably forgot to multiply 120 by -4.

OpenStudy (anonymous):

im confused, could you maybe write down the stps?

OpenStudy (anonymous):

yes you are right c=-480

OpenStudy (mathmate):

@surjithayer wrote down all the steps, the last equation should read \( 11x^2 -28x-480=0 \) from which you can solve for the unknowns using the quadratic formula surjithayer supplied.

OpenStudy (anonymous):

how did you get that out of \[x^2+(1/2x+11)^2=(2x+1)^2\]

OpenStudy (mathmate):

by expansion using the identity on each of the three terms: \((a+b)^2 = a^2 + 2ab + b^2 \) but watch the signs.

OpenStudy (anonymous):

could you explain that more please? im still confused

OpenStudy (anonymous):

correction \[-12^{2}+x ^{2}+28x+480=0\] \[11x ^{2}-28x-480=0\] now you can use the formula given above.

OpenStudy (anonymous):

im so confused. this is what i did. \[x^2+(\frac{ 1 }{ 2 }+11)^2=(2x+1)^2\]\[x^2+\frac{ 1 }{ 2 }x^2+121=4x^2+1\]\[-\frac{ 11 }{ 4 }+120=0\]

OpenStudy (anonymous):

oh sorry \[\frac{ -11 }{ 4}x^2+120=0\]

OpenStudy (anonymous):

@mathmate @surjithayer

OpenStudy (mathmate):

You have missed out the cross terms (2ab) in the identity: \( (a+b)^2=a^2+2ab+b^2 \) If you put those in, you will get exactly the same answer as surjithayer did.

OpenStudy (anonymous):

so how do i get from where i am to your guys's answer

OpenStudy (mathmate):

I think @surjithayer is working on it. It'll be worth the wait!

OpenStudy (anonymous):

\[X=\frac{ +28\pm \sqrt{\left( -28 \right)^{2}-4*11*-480} }{ 2*11 }\] \[x=\frac{ 28\pm \sqrt{784+21120} }{ 22 }\] \[x=\frac{ 28\pm \sqrt{21904} }{22 }=\frac{ 28\pm148 }{22 }=\frac{ 170 }{22 },\frac{ -120 }{ 22 }\] now you can find

OpenStudy (anonymous):

i dont think thats correct, i have an answer key but it just doesnt have the work. the answers are * and -60/11

OpenStudy (anonymous):

@surjithayer

OpenStudy (mathmate):

His work is correct up to \( \frac{28\pm 148}{22} \). He probably meant for you to check his work. Start from there and get your final answer.